If a demand curve is linear, the price elasticity of demand is constant along it.


A. True
B. False
1
Expert's answer
2017-06-01T12:08:10-0400
A= \int V1V2PdV=P∆V=ϑR∆T=1∙8.314∙(373-273.2)=829.7 (J)